OpenStudy (valirium):

Write the expression as a single logarithm. log3 40 - log3 10

OpenStudy (babu):

log3 4

OpenStudy (valirium):

How'd ya get that?

OpenStudy (valirium):

@babu

OpenStudy (babu):

let me write the expression and paste it here, please wait

OpenStudy (babu):

we know, log(base) X - log(base) Y = log(base) (X/Y) so, log3 40 - log3 10 = log3 (40/10) = log3 4 please note that, the base must be same for the expressions involved all the best!

OpenStudy (valirium):

Okay thank you! :)